# Thinking twice inside the box: is Wigner’s friend really quantum?

### v1

@@ III Wigner’s friend in context beyond quantum theory
III.1 Restrictions on descriptions of physical systems and agents
In the previous section, we have argued that the conjunction of the four metaphysical assumptions by Bong et al. [7] are also violated in some classical thought experiments – in other words, quantum theory is not needed to demonstrate that these assumptions cannot always be satisfied. In this section, we will take a more systematic perspective and argue that Wigner’s friend is best understood as a single instance of a larger class of enigmas within physics and philosophy, exceeding the domain of quantum theory. To identify the common thread of these and other conceptual puzzles, it will be helpful to introduce some simple but arguably useful terminology.
Definition 1 (Informal: Restrictions P and A).
Suppose we are given a physical theory (for example quantum theory), perhaps together with a set of plausible additional assumptions (such as certain types of locality or causality assumptions). Then “Restriction P” and/or “Restriction A” may apply, which we define as follows:
Restriction P: For some experiments, the theory cannot provide us with a joint probabilistic description of all relevant properties of all involved physical systems.
@@ III.1 Restrictions on descriptions of physical systems and agents
Importantly, in this formulation, we assume that the restrictions hold even if we are given a complete description of the corresponding experiment within the theory. Note that the assumptions are not meant to be about the physical world, but about the desired probabilistic descriptions.
@@
The term “joint probabilistic description” refers to the following formal structure: a probability space $(\Omega,\mathcal{F},P)$, where $\Omega$ is a sample space, a $\sigma$-algebra $\mathcal{F}$ of events, and a probability measure $P$. We assume that it contains all relevant propositions about properties of physical systems (in the case of Restriction P), or about observations by agents (for Restriction A). Note that Restrictions A and P are possible properties of any given theory and set of assumptions, rather than of a specific implementation.
The paradigmatic example of Restriction P is given by quantum theory’s prediction of the violation of Bell inequalities: assuming Locality Causality and No Superdeterminism, this violation implies that it is impossible to describe, say, the polarisation degrees of freedom of two photons in terms of a collection of two random variables, if their quantum-mechanical description is given by, for example, a singlet state. Thus, Restriction P applies to quantum theory, under the additional assumptions of Locality Causality and No Superdeterminism. Note that these two principles are here not understood as assumptions about the physical world, but about the desired probabilistic description (i.e. what we normally call a “hidden variable” model).
The violation of a Bell inequality is a particularly strong form of Restriction P: not only does quantum theory give us a characterisation of the two polarisations that is more general than any classical probabilistic description, but Bell’s theorem actually predicts the fundamental impossibility to obtain such a description. Hence, every potential future physical theory that predicts a violation of a Bell inequality (as indeed all empirically successful theories now ought to) must also be subject to Restriction P, if we assume Local Causality and No Superdeterminism for the probabilistic descriptions. On the other hand, consider a theory that describes a class of experiments in the regime of classical physics or every-day life, such as classical statistical mechanics. In such case, Restriction P does not apply.
@@
Let us now turn to Restriction A and clarify its meaning by relating it to the original Wigner’s friend thought experiment and their classical variants. Consider one of these experiments, and a theory (say, classical statistical mechanics) that describes it. We can imagine that every agent that is involved in the experiment is initially given a complete description of the experiment. In particular, this will enable the agents to form beliefs about certain facts (such as: what they will see, what the other agents will see, etc.), and such beliefs may be implemented as probability assignments. Accordingly, all agents will start with the same beliefs. However, after some stages of the experiment are concluded, some observers will have learned new facts (obtained some experimental outcomes) that other agents have not. Some of the agents will therefore update their beliefs and some will not – hence the agents will at some point have different beliefs and statistical descriptions. This is to be expected in essentially all statistical theories. The absence of Restriction A does not imply that all agents will always assign the same probabilities: Restriction A is something more dramatic than different agents holding different beliefs.
@@
In particular, if Restriction A applies, then the theory must either fail to describe all statistical empirical observations, or it must be impossible for any external observer (say, a physicist who oversees the experiment) to record all observations of all agents and obtain statistics by repeating the experiment. That is, if the latter is possible, then the external observer can obtain a joint statistical description of all agents’ observations by experiment, and if Restriction A applies, then the theory cannot predict these statistics. This would render the theory empirically incomplete.
This also shows that Restriction A does not apply to situations in quantum physics in which all agents are modelled as physical systems and share a common Heisenberg cut: in this case, an external observer can obtain records of all the other agents’ observations, since they are on the same side of the cut. Thus, repeating the experiment many times, this observer can empirically estimate the probabilities, and the result must be predicted by quantum theory. This is consistent with Vilasini and Woods’ argument that the freedom of choice of Heisenberg cut is the essential element to consider in EWFSs [55].
As a first example of Restriction A, consider Bong et al.’s [7] metaphysical notion of Absoluteness of Observed Events (AOE) (“an observed event is a real single event, and not relative to anything or anyone”):
@@
The negation of AOE is an instance of Restriction A: the prediction of the violation of a Local Friendliness inequality, as demonstrated in [7], proves that Restriction A applies to quantum theory, if supplemented by the assumptions of Locality  [^1] Note that the Locality assumption of [7] is an assumption only about variables that are actually observed (by at least one agent), whereas the locality assumption in Bell’s theorem is about the totality of all (hidden) variables, observed or not. This also explains why it is insufficient to consider simple scenarios such as two observers (say, Alice and Bob) locally observing the outcomes of a Bell experiment violating the CHSH inequality, to conclude that a joint probabilistic description of their observations via local hidden variables cannot exist, and from this to claim that Restriction A applies. This conclusion would rest on a less interesting notion of “Restriction A”, defined for quantum theory with a locality assumption about all the hypothetical observations that Alice and Bob could have made, rather than on their actual observations. and No Superdeterminism. Moreover, it is a particularly strong instance of Restriction A: every other (say, future) theory that makes the same statistical predictions for the EWF scenario of [7] as quantum theory must also be subject to Restriction A, if supplemented by Locality and No Superdeterminism.
@@
To see this, assume Locality and No Superdeterminism, and recall the four-party setup of [7] (the three-party version has been described in Subsection II.1 above). The additional assumption of AOE is then equivalent to the claim that for every fixed choice of settings $x,y$, there is a joint probability distribution $P(a,b,c,d|x,y)$ [7] on the four observations $a,b,c,d$ made by the four agents Alice, Bob, Charlie, and Debbie, and that Alice’s outcome always equals Charlie’s outcome if $x=1$, while Bob’s outcome always equals Debbie’s outcome, if $y=1$. Quantum theory predicts experiments where it at least appears that we can enforce these equality conditions, but observe a violation of a Local Friendliness inequality, implying (under our assumptions) that AOE is false. This implies that there must be at least one pair of settings $x,y$ such that there is no joint probability distribution $P(a,b,c,d|x,y)$ that describes the observations of the four agents. This is an instance of Restriction A.∎
@@
That is, the theoretical results of [7] prove that Restriction A holds for quantum theory. Moreover, the experimental observation of a violation of a Local Friendliness inequality would prove that all empirically adequate future physical theories would also be subject to Restriction A if supplemented by the assumptions of Locality and No Superdeterminism.
@@
This raises the question of what would count as an experimental demonstration. In fact, the violation of a Local Friendliness inequality has been experimentally demonstrated [7], but the role of the “Friend” was played by the path degree of freedom of a single photon. A single photon is clearly not an agent, but what, then, is an agent? What do we mean by “agent” in Restriction A? And while some physical systems clearly do not count as agents, shouldn’t we say that every agent is a physical system? Hence, isn’t Restriction A just a special case of Restriction P?
We take the position that it is not self-evident that we always have to identify agents with physical systems, and that there can be situations where an agent is something else than simply a given material object, or a degree of freedom described by a standalone Hilbert space (as quantum information theorists like to characterise systems). For example, one may hold the view that agents, or persons, are more like patterns than like objects. In the Philosophy of Mind, this would amount to a view where the mind is not the material of my brain, but the associated set of patterns [56]. Moreover, Everettians in particular may want to say that in some cases, more than one person supervenes on a single physical quantum system, and this is one possible reaction to the WF-type thought experiments. In particular, saying that an agent need not be the same as a physical system does not contradict Physical Supervenience: we find many things in our physical world that we can talk about (such as water waves, gene sequences, or computer programs) that are not in direct one-to-one correspondence with standalone physical objects or systems. Thus, not all instances of Restriction A are necessarily special cases of Restriction P.
@@
With this definition in mind, let us consider how Restriction A relates to Thought Experiments 1 and 2:
@@
Restriction A applies to the classical Thought Experiments 1 and 2. In more detail, while classical physics may give us a complete description of the thought experiments, and of Wigner’s observations, it does not give us a probabilistic description of Freya’s observations.
@@
We might be inclined to say that the Friend should be indifferent and assign probability $1/2$ to each of the two alternatives. Indeed, Elga has defended a “Principle of Indifference” [40], claiming that one should assign uniform probabilities in cases of self-locating uncertainty. We describe Elga’s principle and its relation to our thought experiments in Appendix V.4. However, whatever probability value $p$ the Friend decides to put here, this number cannot come from classical physics, but it amounts to an additional choice that has to be made. This can also be seen by noting that it is impossible to verify this probability assignment empirically: external observers (such as Wigner, and potentially other physicists, in particular those that write publications) cannot repeat the experiment many times and obtain empirical statistics on the number of runs in which Freya actually saw green versus blue (moreover, we have argued in Section II.2 that the question itself has no straightforward meaning, since the referent is ambiguous, or even multivalued, if described from an external perspective). This is in stark contrast to probability assignments in, say, classical statistical mechanics: all mechanical properties can be measured, statistics can be obtained externally, and thus all probability assignments can in principle be tested empirically [^2] Here we have assumed that the colour (either green or blue) that Freya will find herself seeing, which we have imagined being described by a probability distribution $(p,1-p)$, is uncorrelated with all facts of the world accessible to external observers. It can be argued that this claim is implicit in the assumption that our scenario involves duplication: both resulting versions of Freya (having seen green or blue, respectively) should look like identically legitimate successors of initial Freya for every external observer. For example, both copies should answer “Freya!” when asked for their names; similarly, no other facts of the world should have anything to say as to which one of the two is “more likely initial Freya’s successor” than the other one. In principle, we could imagine that this assumption is not true: perhaps the value of an external variable (say, a classical bit somewhere in thermal radiation) predetermined whether Freya will see green or blue, rendering the other copy’s impression that she has had an earlier life as “Freya before the experiment” an illusion. But any theory that would make such a prediction is different from, or at least more than, classical mechanics, classical statistical mechanics, or any other theory we have previously formulated.. Hence, the value of $p$ must lie beyond the scope of classical physics. We would need to use “classical physics plus Elga’s Principle of Indifference” [40], for example, in order to avoid Restriction A.
Since Thought Experiment 1 was explicitly constructed as an analogue for Wigner’s friend, the following claim might come as a surprise:
@@
Restriction A does not apply to the original WF experiment (one Wigner, one Friend) for quantum theory.
More specifically, Restriction A does not apply to any interpretation of quantum mechanics that postulates that is is meaningful and correct for the Friend to use the Born rule, despite the fact that a naive use of the collapse postulate could lead to incorrect predictions of Wigner’s outcome for a measurement in the entangled basis. That is, for such interpretations, both Wigner and his Friend can ask questions such as “What will I see next?”, and quantum theory can assign probabilities to the outcomes via the Born rule – in disanalogy with Thought Experiment 1, for which classical physics by itself cannot assign probabilities to Freya’s subjective outcomes.
On the other hand, Thought Experiment 3 is an even more interesting case of Restriction A:
@@
Restriction A applies to Thought Experiment 3 as described by classical physics. Moreover, even if we supplement classical physics with any probability rule whatsoever (not necessarily the Principle of Indifference), which informs Freya about what she should believe about her future observations in a way that is not completely ignoring her subsequent multiplicity $M$, then the resulting theory will be subject to Restriction A.
In the description of the experiment, we have implicitly assumed the exchangeability of all labs: that permuting the order of all laboratories would not change any of the outcome statistics, as experienced by Freya, Wigner, or any other agent. Now, if we have $N$ laboratories, let us denote Freya ending up in laboratory $j$, and in a world where the coin toss in this lab has produced Heads, by $P_{F}(H,j)$. Then
$P_{F}(H)=\sum_{j=1}^{N}P_{F}(H,j)=\sum_{j=1}^{N}P_{F}(H|j)P_{F}(f),$
and assuming via exchangeability that all $P_{F}(j)$ are identical (and that so are the $P_{F}(H|j)$), we get $P_{F}(H)=P_{F}(H|j=1)$, and we can restrict our attention to the case of $N=1$ laboratories.
Suppose Restriction A does not apply, and consider a joint probability space describing the observations of all agents. Then the event “Wigner observes a Heads outcome” happens if and only if the event “Freya observes a Heads outcome” happens, no matter which of the possible copies Freya experiences herself ending up (whatever that means). But the probability of the former ($1/2$) is independent of the number $M$ of potential copies of Freya, and hence so must be the probability of the latter. ∎
Hence, if Freya is convinced that the number of multiplicities $M$ should have something to tell her about the outcome she should expect to see, then she cannot consult her physics textbooks to tell her what probability she should assign. No joint probability space could describe her and Wigner’s assignments, and the situation resembles others that have been argued to point at an agent-dependence of facts [54, 57, 58, 59, 60, 61, 62]. Moreover, if we interpret Wigner’s “halfer” and Freya’s “thirder” assignments as private but objective chances in some sense, then we could regard Freya and Wigner as probabilistic zombies relative to each other, as described in [63], resembling speculations of [64].
@@
Whatever resources beyond the physics textbooks Freya consults to obtain a probability assignment that is not independent of $M$ (such as Elga’s Principle of Indifference), this will block her ability to model her observations together with those of other agents (in particular, of Wigner) in terms of a joint probability space, as usually dictated by the postulates of probability theory. In particular, this will block her ability to “pool” her knowledge with that of other agents, a phenomenon that has prominently been derived within quantum theory in the Frauchiger-Renner thought experiment [5]. The inability to do so raises further questions, for example a question posed by Renner in the context of quantum theory [65]: Given that Restriction A blocks several instances of reasoning that combines the knowledge or belief of several agents, what would consitute useful sufficient conditions for when such reasoning is still possible? For a suggestion of how to address this question in the quantum case, see e.g. the work by Renes [66].
Restriction A from the perspective of personhood
@@
In Section II, we have argued for the relevance of philosophical positions on personal identity to make sense of the thought experiments. In this subsection, we have so far circumvented this conclusion by relying on a pragmatic definition of “agent”, see Definition 2 above. While this is sufficient to introduce the notion of Restriction A and discuss its applicability to different physical theories and scenarios, it is still interesting to ask how Restriction A would be interpreted or understood by proponents of these different philosophical views on personhood, as summarised in Appendix V.2.
@@
A possible response by a supporter of Parfit’s views could be as follows. They might say that the attempt to regard the “agent” in Restriction A as an ontological existing entity, a person, would be misguided and would render this definition meaningless. For example, in Thought Experiment 2, there is no ontological notion of “Freya” as the person that entered the experiment and that also saw either specific colour. In fact, they might interpret the appearance of Restriction A as a symptom of the misguided attempt to reason about such a hypothetical entity when there is, as a matter of fact (of the world), actually none, reflected in the structural observation that the world’s probability space does not carry a corresponding random variable that persists over the different temporal stages of the experiment. However, they might still agree that Restriction A, together with our pragmatic definition of agent, is a meaningful notion: Parfit wrote that “being destroyed and replicated is about as good as ordinary survival” [29] — and then, arguably so is the situation involved in a WF-type or duplication experiment. And since it is rational under ordinary survival to attempt to construct theories that help us to assign probabilities to our possible future observations, a Parfitian should then not deny that agents are rational who attempt the same when facing WF-type situations, whatever the word “our” means in the description of this endeavour.
[…]

### v2

@@ III Wigner’s Friend in context beyond quantum theory
In this section, we offer a novel way of interpreting the feature at the core of EWF scenarios – in particular, something structural in nature, that may also apply to other physical theories, e.g. classical ones. In Subsection III.1, we define the concept of Restriction 1, show how it applies to the previously described scenarios, and relate it to the existing, metaphysical claims concerning absoluteness [7] or epistemic horizons [56]. In Subsection III.2, we review how Bell inequality violations in quantum theory can always be lifted to instances of Restriction 1, and comment on the relation to Fine’s Theorem and the relation between agents and physical systems. Finally, in Subsection III.3, we discuss the seemingly unrelated Boltzmann brain problem of cosmology in terms of Restriction 1.
III.1 Definition of Restriction 1
In the previous section, we have argued that important features of Wigner’s Friend-type scenarios have analogues in classical duplication experiments. In this section, we will take a more systematic perspective and identify a common structural element of these and other conceptual puzzles, which we define as follows:
Definition 1.
Suppose we are given a physical theory (for example quantum theory), perhaps together with a set of plausible additional assumptions (such as certain types of locality or causality assumptions). Then “Restriction 1” may apply, which we define as follows:
@@ III.1 Definition of Restriction 1
Importantly, in this formulation, we assume that the restrictions hold even if we are given a complete description of the corresponding experiment within the theory.
@@
The term “joint probabilistic description” refers to the following formal structure: a probability space $(\Omega,\mathcal{F},P)$, where $\Omega$ is a sample space, a $\sigma$-algebra $\mathcal{F}$ of events, and a probability measure $P$, containing the observations as random variables. Note that Restrictions 1 is a possible property of a given theory and set of assumptions, rather than of a specific implementation.
Restriction 1 can apply to $N$ agents, and the cases $N=1$ and $N\geq 2$ have very different interpretations: Restriction 1 for $N=1$ means that the respective agent cannot use the theory to obtain a probabilistic prediction for its future observations; Restriction 1 for $N\geq 2$ means that such a prediction may be obtained, but not by predicting the observations of all agents and considering the marginal distribution for one of them. In some sense, the observations of the agents do not fit into a joint probability space of shared facts.
@@
Let us first understand a necessary condition for Restriction 1 to apply. To this end, consider an experiment involving several agents, and a theory (say, classical statistical mechanics) that describes it. We can imagine that every agent that is involved in the experiment is initially given a complete description of the experiment. In particular, this will enable the agents to form beliefs about certain facts (such as: what they will see, what the other agents will see, etc.), and such beliefs may be implemented as probability assignments. Accordingly, all agents will start with the same beliefs. However, after some stages of the experiment are concluded, some observers will have learned new facts (obtained some experimental outcomes) that other agents have not. Some of the agents will therefore update their beliefs and some will not – hence the agents will at some point have different beliefs and statistical descriptions. This is to be expected in essentially all statistical theories. The absence of Restriction 1 does not imply that all agents will always assign the same probabilities: Restriction 1 is something more dramatic than different agents holding different beliefs.
@@
In particular, if Restriction 1 applies, then the theory must either fail to describe all statistical empirical observations, or it must be impossible for any external observer (say, a physicist who oversees the experiment) to record all observations of all agents and obtain statistics by repeating the experiment. That is, if the latter is possible, then the external observer can obtain a joint statistical description of all agents’ observations by experiment, and if Restriction 1 applies, then the theory cannot predict these statistics. This would render the theory empirically incomplete. This leads us to the following claim:
@@
In an empirically complete theory, Restriction 1 only applies in situations where the observations of all agents cannot be externally jointly assessed — that is, in situations (to borrow the language of [56]) that feature an “epistemic horizon”.
That is, an essential ingredient for the demonstration of Restriction 1 is something that plays the role of a “black box” in the experiment. For EWF scenarios, this is the closed laboratory and subsequent unitary reversal of the Friend’s measurement process, which forbids the outcome of their measurement being known intersubjectively. Meanwhile, in classical Thought Experiments II.2 and II.3, the duplication process makes it impossible to even define from an external point of view what “Freya’s outcome” has been, even though Freya may privately be justified in wondering before the experiment what she will observe. In this sense, the random experiment becomes a private experiment, and the epistemic horizon is due to the fact that its outcome is only known to an agent who performs the experiment themself. Similarly, in the Boltzmann brain (BB) problem (to be discussed further in Subsection III.3), the question “am I a BB?” may reveal its answer to the questioner, but not to any external observer since the answer does not correspond to any fact of the world.
This also shows that Restriction 1 does not apply to situations in quantum physics in which all agents are modelled as physical systems and share a common Heisenberg cut: in this case, an external observer can obtain records of all the other agents’ observations, since they are on the same side of the cut. Thus, repeating the experiment many times, this observer can empirically estimate the probabilities, and the result must be predicted by quantum theory. This is consistent with Vilasini and Woods’ argument that the freedom of choice of Heisenberg cut is an essential element to consider in EWFSs [57].
As a first example of Restriction 1, consider Bong et al.’s [7] notion of Absoluteness of Observed Events (AOE) (“an observed event is a real single event, and not relative to anything or anyone”):
Claim 2.
@@
The negation of AOE is an instance of Restriction 1: the prediction of the violation of a Local Friendliness inequality, as demonstrated in [7], proves that Restriction 1 applies to quantum theory, if supplemented by the assumptions of Locality  [^1] Note that the Locality assumption of [7] is an assumption only about variables that are actually observed (by at least one agent), whereas the locality assumption in Bell’s theorem is about the totality of all (hidden) variables, observed or not. This also explains why it is insufficient to consider simple scenarios such as two observers (say, Alice and Bob) locally observing the outcomes of a Bell experiment violating the CHSH inequality, to conclude that a joint probabilistic description of their observations via local hidden variables cannot exist, and from this to claim that Restriction 1 applies. This conclusion would rest on a less interesting notion of “Restriction 1”, defined for quantum theory with a locality assumption about all the hypothetical observations that Alice and Bob could have made, rather than on their actual observations. and No Superdeterminism. Moreover, it is a particularly strong instance of Restriction 1: every other (say, future) theory that makes the same statistical predictions for the EWF scenario of [7] as quantum theory must also be subject to Restriction 1, if supplemented by Locality and No Superdeterminism.
@@
To see this, assume Locality and No Superdeterminism, and recall the four-party setup of [7] (the three-party version has been described in Subsection II.1 above). The additional assumption of AOE is then equivalent to the claim that for every fixed choice of settings $x,y$, there is a joint probability distribution $P(a,b,c,d|x,y)$ [7] on the four observations $a,b,c,d$ made by the four agents Alice, Bob, Charlie, and Debbie, and that Alice’s outcome always equals Charlie’s outcome if $x=1$, while Bob’s outcome always equals Debbie’s outcome, if $y=1$. Quantum theory predicts experiments where it at least appears that we can enforce these equality conditions, but observe a violation of a Local Friendliness inequality, implying (under our assumptions) that AOE is false. This implies that there must be at least one pair of settings $x,y$ such that there is no joint probability distribution $P(a,b,c,d|x,y)$ that describes the observations of the four agents. This is an instance of Restriction 1.∎
@@
That is, the theoretical results of [7] prove that Restriction 1 holds for quantum theory. Moreover, the experimental observation of a violation of a Local Friendliness inequality would prove that all empirically adequate future physical theories would also be subject to Restriction 1 if supplemented by the assumptions of Locality and No Superdeterminism.
@@
This raises the question of what would count as an experimental demonstration. In fact, the violation of a Local Friendliness inequality has been experimentally demonstrated [7], but the role of the “Friend” was played by the path degree of freedom of a single photon. A single photon is clearly not an agent, but what, then, is an agent?
@@
With this definition in mind, let us consider how Restriction 1 relates to Thought Experiment II.2:
@@
Restriction 1 applies to the classical Thought Experiment II.2. In more detail, while classical physics may give us a complete description of the thought experiment, and of Wigner’s observations, it does not give us a probabilistic description of Freya’s observations.
@@
We might be inclined to say that the Friend should be indifferent and assign probability $1/2$ to each of the two alternatives. Indeed, Elga has defended a “Principle of Indifference” [41], claiming that one should assign uniform probabilities in cases of self-locating uncertainty. We describe Elga’s principle and its relation to our thought experiments in Appendix V.4. However, whatever probability value $p$ the Friend decides to put here, this number cannot come from classical physics, but it amounts to an additional choice that has to be made. This can also be seen by noting that it is impossible to verify this probability assignment empirically: external observers (such as Wigner, and potentially other physicists, in particular those that write publications) cannot repeat the experiment many times and obtain empirical statistics on the number of runs in which Freya “actually saw” green versus blue (moreover, we have argued in Section II.2 that the question itself has no straightforward meaning, since the referent is ambiguous, or even multivalued, if described from an external perspective). This is in stark contrast to probability assignments in, say, classical statistical mechanics: all mechanical properties can be measured, statistics can be obtained externally, and thus all probability assignments can in principle be tested empirically [^2] Here we have assumed that the colour (either green or blue) that Freya will find herself seeing, which we have imagined being described by a probability distribution $(p,1-p)$, is uncorrelated with all facts of the world accessible to external observers. It can be argued that this claim is implicit in the assumption that our scenario involves duplication: both resulting versions of Freya (having seen green or blue, respectively) should look like identically legitimate successors of initial Freya for every external observer. For example, both copies should answer “Freya!” when asked for their names; similarly, no other facts of the world should have anything to say as to which one of the two is “more likely initial Freya’s successor” than the other one. In principle, we could imagine that this assumption is not true: perhaps the value of an external variable (say, a classical bit somewhere in thermal radiation) predetermined whether Freya will see green or blue, rendering the other copy’s impression that she has had an earlier life as “Freya before the experiment” an illusion. But any theory that would make such a prediction is different from, or at least more than, classical mechanics, classical statistical mechanics, or any other theory we have previously formulated.. Hence, the value of $p$ must lie beyond the scope of classical physics. We would need to use “classical physics plus Elga’s Principle of Indifference” [41], for example, in order to avoid Restriction 1.
Since aspects of Thought Experiment II.2 were explicitly constructed as an analogue for Wigner’s Friend, it is interesting to compare the two:
@@
Restriction 1 does not apply to the original WF experiment (one Wigner, one Friend). However, if the scenario is extended to allow Wigner to unitarily reverse the Friend’s measurement, and to perform a measurement that is incompatible with the Friend’s, then Restriction 1 applies to these $N=2$ agents. But despite the strong analogy with classical Thought Experiment II.2, Restriction 1 does not apply individually to $N=1$ agent (say, the Friend).
The original WF scenario points out a tension between Wigner’s and the Friend’s descriptions of the quantum state after the Friend’s measurement, but it does not preclude the possibility that the Friend tells Wigner its result after the experiment. Since Wigner and his Friend can (by at least some interpretations) be thought of as sharing a common Heisenberg cut after the Friend’s experiment, it is possible to describe all observations made by both by a joint probability distribution, as provided by quantum theory.
However, the situation is different if measurements in incompatible bases are involved (see e.g. [58, Section 4.1]. It is the possibility of unitarily reversing the Friend’s lab (c.f. [4], or in the terminology of [27], “Wigner’s Enemy”) that, under certain assumptions, precludes the existence of a single, consistent probability table – in particular, because this gives rise to an epistemic horizon [56] (see Claim 1). The scenario with unitary reversal is analogous to our classical Thought Experiment II.2 – but, in the latter, Restriction 1 even applies to Freya individually: classical physics does not give her a probability for what to see in the experiment, regardless of the question of whether this probability would fit into a joint distribution with a superobserver. Quantum theory, on the other hand, does tell the Friend what to expect to observe before the unitary reversal has been implemented: namely, via the Born rule.
Thought Experiment II.3 is an even more interesting case of Restriction 1:
@@
Restriction 1 applies to Wigner and Freya individually in Thought Experiment II.3 as described by classical physics. Moreover, if we supplement classical physics with Elga’s Principle of Indifference and (NUI) to inform Freya and Wigner’s probabilities, then there is no joint probabilistic description for the observations of these two agents.
Clearly, classical physics alone does not provide Freya or Wigner with probabilistic predictions for the coin toss results that they will observe: “Freya (or Wigner) sees Heads” is not even a well-defined event on the probability space of the $N$ coin tosses. Hence, Restriction 1 applies to Freya and to Wigner individually.
Now suppose we supplement classical physics with a probability rule, such as Elga’s Principle of Indifference (or, in fact, any rule for which Wigner’s but not Freya’s probability is independent of $M$), which defines $P_{F}(H)$ and $P_{W}(H)$, and we assume (NUI). As in the proof of Theorem 1, Assumption (NUI) implies that $P_{W}(T)$ and $P_{F}(T)$ are independent of $N$. Hence, consider the case $N=1$. Suppose that Restriction 1 does not hold, then we have a joint probability space with a single distribution $P$ and two events $F_{T}$ and $W_{T}$, describing that Freya and Wigner, respectively, observe Tails. By assumption, since $N=1$, $F_{T}$ happens if and only if $W_{T}$ happens, hence $P_{W}(T)=P(W_{T})=P(F_{T})=P_{F}(T)$. But then, if $P_{W}(T)$ is independent of $M$, so must $P_{F}(T)$. But we have assumed that this is not the case – therefore, Restriction 1 must apply.
The application of Elga’s Principle of Indifference introduces an observer-dependence into the scenario that can be made explicit by phrasing the question in different words: “Will I find myself in a world where the coin shows Heads?”. This formulation makes it appear less paradoxial that the transitivity of knowledge that is assumed by (C) and (CP) is no longer relevant; although Wigner may reason “I am certain that Freya, upon reasoning within the same theory as the one I am using, and having the same exact knowledge of the world as I, is pretty sure that she will find herself in a world where $c=H$ at time $t$”, this is not to say that Wigner should conclude “I am pretty sure that I will find myself in a world where $c=H$ at time $t$”. This situation resembles others that have been argued to point at an agent-dependence of facts [55, 59, 60, 61, 62, 63, 64]. Moreover, if we interpret Wigner’s “halfer” and Freya’s “thirder” assignments as private but objective chances in some sense, then we could regard Freya and Wigner as probabilistic zombies relative to each other, as described in [65], resembling speculations of [66].
@@
Whatever resources beyond the physics textbooks Freya consults to obtain a probability assignment that depends on $M$ (such as Elga’s Principle of Indifference), this will block her ability to model her observations together with those of other agents (in particular, of Wigner) in terms of a joint probability space, as usually dictated by the postulates of probability theory. In particular, this will block her ability to “pool” her knowledge with that of other agents, a phenomenon that has prominently been derived within quantum theory in the Frauchiger-Renner thought experiment [5]. The inability to do so raises further questions, for example a question posed by Renner in the context of quantum theory [67]: Given that Restriction 1 blocks several instances of reasoning that combines the knowledge or belief of several agents, what would consitute useful sufficient conditions for when such reasoning is still possible? For a suggestion of how to address this question in the quantum case, see e.g. the work by Renes [68] or Vilasini and Woods [57].
Restriction 1 from the perspective of personhood
@@
In Section II, we have argued for the relevance of philosophical positions on personal identity to make sense of the thought experiments. In this subsection, we have so far circumvented this conclusion by relying on a pragmatic definition of “agent”, see Definition 2 above. While this is sufficient to introduce the notion of Restriction 1 and discuss its applicability to different physical theories and scenarios, it is still interesting to ask how Restriction 1 would be interpreted or understood by proponents of these different philosophical views on personhood, as summarised in Appendix V.2.
@@
A possible response by a supporter of Parfit’s views could be as follows. They might say that the attempt to regard the “agent” in Restriction 1 as an ontological existing entity, a person, would be misguided and would render this definition meaningless. For example, in Thought Experiment II.2, there is no ontological notion of “Freya” as the person that entered the experiment and that also saw either specific colour. In fact, they might interpret the appearance of Restriction 1 as a symptom of the misguided attempt to reason about such a hypothetical entity when there is, as a matter of fact (of the world), actually none, reflected in the structural observation that the world’s probability space does not carry a corresponding random variable that persists over the different temporal stages of the experiment. However, they might still agree that Restriction 1, together with our pragmatic definition of agent, is a meaningful notion: Parfit wrote that “being destroyed and replicated is about as good as ordinary survival” [30] — and then, arguably so is the situation involved in a WF-type or duplication experiment. And since it is rational under ordinary survival to attempt to construct theories that help us to assign probabilities to our possible future observations, a Parfitian should then not deny that agents are rational who attempt the same when facing WF-type situations, whatever the word “our” means in the description of this endeavour.
Sider’s ‘Stage view’ might say something similar. Whilst Sider’s account of personhood is not as metaphysically reductionist as Parfit’s, he too considers the understanding of Freya as a single, persisting person to be mistaken. His own ontology, consisting of three-dimensional person stages, would see there as being multiple persons existing over the course of the experiment. Statements that refer to future or past persons may therefore need to be relativised to a particular class of person stages. Restriction 1 may then be viewed as a consequence of mistakenly taking the aggregates of person stages to be fundamental (c.f. Lewisians, see next paragraph), rather than the stages themselves. In particular, if we consider just the person stage Freya0, prior to the duplication process of Thought Experiment II.2, it is perhaps unsurprising that our physical theories do not offer probabilities for whether Freya0 will become FreyaB or FreyaG – because this question presupposes mistaken notions of persisting persons beyond simple similarity. Equally, for EWF scenarios, perhaps the contradiction is also rooted in an implicit ontology of persons, applied to scenarios for which the identification and composition of person stages is non-trivial. Nevertheless, as with Parfit’s view, FreyaB or FreyaG are person-stages that should matter to Freya0, due to their connection via an “I-relation” – therefore it still seems natural to ask questions about, essentially, how much they should matter.